# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ Abstract
These are self-contained lecture notes for spectral independence. For an $n$-vertex graph, the spectral independence condition is a bound on the maximum eigenvalue of the $n\times n$ influence matrix whose entries capture the influence between pairs of vertices, it is closely related to the covariance matrix. We will present recent results showing that spectral independence implies the mixing time of the Glauber dynamics is polynomial (where the degree of the polynomial depends on certain parameters). The proof utilizes local-to-global theorems which we will detail in these notes. Finally, we will present more recent results showing that spectral independence implies an optimal bound on the relaxation time (inverse spectral gap) and with some additional conditions implies an optimal mixing time bound of $O(n\log{n})$ for the Glauber dynamics.
We also present the results of Anari, Liu, Oveis Gharan, and Vinzant (2019) for generating a random basis of a matroid. The analysis of the associated bases-exchange walk utilizes the local-to-global theorems used for spectral independence with the Trickle-Down Theorem of Oppenheim (2018) to analyze the local walks. Our focus in these notes is on the analysis of the spectral gap of the associated Markov chains from a functional analysis perspective, and we present proofs of the associated local-to-global theorems from this same Markov chain perspective.
@@ 1 Introduction
These are a comprehensive set of lecture notes on spectral independence and its implications on the mixing time of the single-site update Markov chain known as the Glauber dynamics.
Spectral independence is a simple, yet seemingly quite powerful technique to obtain optimal upper bounds on the convergence rate of Markov chains to its stationary distribution. In these notes we will introduce the spectral independence technique, explain the basic properties of the associated influence matrix, and discuss connections to simplicial complexes. We then prove or at least provide the main details for the main results that spectral independence implies fast mixing of the Glauber dynamics.
We also present the prominent result of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19] which shows fast mixing of a Markov chain, known as the bases-exchange walk, for randomly sampling bases of a matroid. This utilizes the local-to-global theorems presented for spectral independence and the Trickle-Down Theorem of Oppenheim [Opp18] to bound the spectral gap of the local walks.
[…]
@@ 1.1 Setting
These notes study high-dimensional distributions defined by a graph, this corresponds to spin systems in statistical physics and undirected graphical models in machine learning. Let $G=(V,E)$ be a graph on $n=|V|$ vertices.
Consider a distribution $\mu$ on a subset of $\{0,1\}^{V}$. Let
$\Omega=\{\sigma\in\{0,1\}^{V}:\mu(\sigma)>0\}.$
All of the results in these notes generalize to distributions defined on $\{1,\dots,q\}^{V}$ for integer $q\geq 2$; we will comment on these generalizations later in these notes.
An example to keep in mind is the so-called hard-core model which we define next.
@@ 1.2 Running example: Hard-core Model
The input to the hard-core model is a graph $G=(V,E)$ and an activity $\lambda>0$. The hard-core model is an idealized model of a gas where $\lambda$ corresponds to the fugacity of the gas. Configurations of the model are defined on independent sets of $G$; recall, an independent set is a subset $S$ of vertices which does not contain an edge, i.e., for all $\{x,y\}\in E$ either $x\notin S$ and/or $y\notin S$. Let $\Omega$ denote the collection of independent sets of $G$ (regardless of their sizes).
For an independent set $\sigma\in\Omega$, we can view $\sigma$ as an $n$-dimensional vector in $\{0,1\}^{n}$ where the $i$-th coordinate is assigned 1 if the $i$-th vertex is in $\sigma$ and is assigned 0 otherwise. Each independent set $\sigma\in\Omega$ is assigned a weight
@@
where $|\sigma|$ is the number of vertices in the independent set $\sigma$. The Gibbs distribution $\mu=\mu_{G,\lambda}$ is the probability distribution proportional to the weights. For $\sigma\in\Omega$,
@@
where the partition function $Z=Z_{G,\lambda}$ is the normalizing factor $Z=\sum_{\eta\in\Omega}\lambda^{|\eta|}$.
1.3 Glauber dynamics
The Glauber dynamics, also known as the Gibbs sampler, is a simple Markov chain designed to sample from the Gibbs distribution. We begin with the definition of the Glauber dynamics for the special case of the hard-core model (moreover one can further consider the case $\lambda=1$ for simplicity).
@@ 1.3 Glauber dynamics
    If $X^{\prime}\in\Omega$ (i.e., it is a valid independent set) then set $X_{t+1}=X_{t}$ and otherwise set $X_{t+1}=X_{t}$.
Notice that the Glauber dynamics is irreducible (since all states can reach the empty set) and aperiodic (since there is a self-loop of not changing the state at $v$) and hence the chain is ergodic where the unique stationary distribution is the Gibbs distribution.
We can now present the general definition of the Glauber dynamics for more general distributions. For a distribution $\mu$ on $\Omega\subset\{0,1\}^{V}$, the Glauber dynamics is defined as follows. From a state $X_{t}\in\{0,1\}^{V}$,
@@
    Choose $X_{t+1}(v)$ from the conditional Gibbs distribution $\mu(\sigma(v)|\sigma(w)=X_{t}(w)\mbox{ for all }w\neq v\})$, i.e., fix the spin at all vertices except $v$ and resample the spin/label at $v$ conditional on the fixed configuration on the rest of the vertices.
@@
The mixing time of the Glauber dynamics, denoted as $T_{\mathrm{mix}}(\varepsilon)$ for $\varepsilon>0$, is the minimum number of steps $t$ such that, for the worst initial state $X_{0}$, the distribution of $X_{t}$ is within total variation distance $\leq\varepsilon$ of the stationary distribution $\pi$. We will refer to $T_{\mathrm{mix}}=T_{\mathrm{mix}}(1/4)$ as the mixing time. The choice of constant $1/4$ is somewhat arbitrary as one can “boost” the mixing time to obtain, for any $\varepsilon>0$, $T_{\mathrm{mix}}(\varepsilon)\leq T_{\mathrm{mix}}(1/4)\times\lceil\log_{2}(1/\varepsilon)\rceil$. Oftentimes, the constant $1/2e$ is used in place of $1/4$ in the definition of the mixing time and then the corresponding boosting has $\ln(1/\varepsilon)$ in place of $\log_{2}(1/\varepsilon)$, see [LP17, Section 4.5].
@@
We say the mixing time is optimal if $T_{\mathrm{mix}}=O(n\log{n})$ as this matches the lower bound established by Hayes and Sinclair [HS07] for any graph of constant maximum degree $\Delta$. The proofs of optimal mixing in these notes will yield an $O(n\log(n/\varepsilon))$ mixing time bound for any $\varepsilon>0$, which is better than obtained from the standard boosting technique.
@@
In contrast to the mixing time, a weaker notion of convergence is the relaxation time. Let $\lambda_{1}=1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-1$ denote the eigenvalues of the Glauber dynamics. Note, $\lambda_{1}$ corresponds to the principal eigenvector which is the stationary distribution $\pi$. Moreover, $\lambda_{2}<1$ since the Glauber dynamics is an ergodic Markov chain and hence the stationary distribution is unique. In addition, $\lambda_{N}>-1$ since the Markov chain is aperiodic. The relaxation time is defined as $T_{\mathrm{relax}}:=(1-\lambda_{*})^{-1}$ where $\lambda_{*}=\max\{\lambda_{2},|\lambda_{N}|\}$. The relaxation time measures the decay rate of the variance of the chain with respect to the stationary distribution, see Section 3 for a more detailed discussion.
The relaxation time is often the critical quantity for approximate counting algorithms in which the aim is to estimate the partition function $Z$ using random samples from the Gibbs distribution. In particular, an optimal mixing time of $O(n\log(n/\varepsilon))$ for the Glauber dynamics implies an $O((n/\delta)^{2}\log(n)\log(n/\delta))$ time algorithm for estimating the partition function $Z$ within an $(1\pm\delta)$-multiplicative factor, for any $\delta>0$, see [ŠVV09, Hub15, Kol18].
1.4 Spectral Independence
@@
Spectral independence was introduced by Anari, Liu, and Oveis Gharan [ALO20]. It is defined by a $n\times n$ influence matrix which captures the pairwise influence or correlations between pairs of vertices.
@@
Let $G=(V,E)$ be a graph where $V=\{1,\dots,n\}$, and $\mu$ be a distribution on a subset of $\{0,1\}^{V}$. Let $\Psi$ be the following real-valued $n\times n$ matrix. For $1\leq i,j\leq n$,
$\Psi(i\rightarrow j)=\Psi(i,j):=\mu\left[\sigma(j)=1\mid\sigma(i)=1\right]-\mu\left[\sigma(j)=1\mid\sigma(i)=0\right].$
We will need to extend the above definition to arbitrary “pinnings” which are a fixed assignment $\tau$ to a subset $S$ of vertices. For $S\subset[n]=\{1,\dots,n\}$, a pinning on $S$ is an assignment $\tau:S\rightarrow\{0,1\}$. Let $\mathcal{P}$ denote the collection of valid pinnings where we assume $\tau\in\mathcal{P}$ is valid in the sense that there exists $\sigma\in\Omega$ where $\sigma(S)=\tau$; see Section 4.5 for a more detailed definition. Throughout these notes we use the notation $\sigma(S)$ to refer to the restriction of $\sigma$ to the set $S$.
@@
For a pinning $\tau\in\mathcal{P}$, let $\mu_{\tau}$ denote the conditional Gibbs distribution, i.e., the distribution $\mu$ conditional on the fixed assignment $\tau$ on $S$. For $\sigma\in\Omega$, we have
$\mu_{\tau}(\sigma)=\mu(\sigma\mid\sigma(S)=\tau)=\frac{\mu(\sigma(\overline{S})\cup\tau(S))}{\mu(\tau)},$
where
$\mu(\tau)=\sum_{\eta\in\Omega:\atop\tau(S)=\eta(S)}\mu(\eta).$
@@
For a pinning $\tau\in\mathcal{P}$ where $\tau:S\rightarrow\{0,1\}$ for $S\subset V$, let $T\subset V\setminus S$ denote the set of vertices which are “free” in the following sense:
$i\in T\iff\mu_{\tau}(\sigma(i)=1)>0\mbox{ and }\mu_{\tau}(\sigma(i)=0)>0.$
@@
Note, a vertex $i\in V\setminus(S\cup T)$ is “frozen”, i.e., it can only attain one spin and hence we can fix the configuration on that vertex and remove the corresponding two rows/columns from the matrix $\Psi_{\tau}$.
@@
$\Psi_{\tau}(i\rightarrow j)$ $:=\mu_{\tau}\left[\sigma(j)=1\mid\sigma(i)=1\right]-\mu_{\tau}\left[\sigma(j)=1\mid\sigma(i)=0\right]$ (1)
$=\mu\left[\sigma(j)=1\mid\sigma(i)=1,\sigma(S)=\tau\right]-\mu\left[\sigma(j)=1\mid\sigma(i)=0,\sigma(S)=\tau\right].$
Note the empty pinning $\emptyset$ is always a valid pinning and hence $\Psi=\Psi_{\emptyset}$.
The matrix $\Psi$ can be asymmetric and the entries of $\Psi$ can be positive or negative. Nevertheless all of its eigenvalues are real, see Lemma 2.1. Intuitively the fact that $\Psi$ is positive semidefinite is because it is closely related to the covariance matrix $\mathrm{Cov}$. In particular, the influence matrix $\Psi$ is equal to a positive diagonal matrix times the covariance matrix $\mathrm{Cov}$ and then the fact follows from the fact that the covariance matrix is positive semidefinite.
Since all eigenvalues of $\Psi$ are real numbers we can denote the maximum eigenvalue by $\lambda_{\max}(\Psi)$. Now we can define spectral independence.
@@
For $\eta>0$, we say that $\mu$ is $\eta$-spectrally independent if for all pinnings $\tau\in\mathcal{P}$, $\lambda_{\max}(\Psi_{\tau})\leq 1+\eta$.
Note that the spectral independence condition only depends on the distribution $\mu$, it has no explicit dependence on the Glauber dynamics. The definition was extended to non-binary spin spaces, such as the Potts model and colorings, in [FGYZ21, CGŠV21] (see also [CLV21, ${}^{+}$22] for a general formulation).
When $\mu$ is a product distribution then $\eta=0$. Our goal is to show that $\eta$ is constant.

### v4

# Spectral Independence and Local-to-Global Techniques for Optimal Mixing of Markov Chains

@@ Abstract
This monograph is an exposition on an exciting new technique known as spectral independence, which has been instrumental in analyzing the convergence rate of Markov Chain Monte Carlo (MCMC) algorithms. For a high-dimensional distribution defined on labelings of the vertices of an $n$-vertex graph, the spectral independence condition, introduced by Anari, Liu, and Oveis Gharan (2020), is a bound on the maximum eigenvalue of the $n\times n$ influence matrix whose entries capture the influence between pairs of vertices (closely related to the covariance between the variables). In the first part of the monograph, we present results showing that spectral independence (and related techniques) imply fast mixing of simple Markov chains such as the Gibbs sampler. These proofs utilize local-to-global theorems which we will detail in this work.
We focus on two applications: the hard-core model on independent sets of a graph (which is a binary graphical model) and random bases of a matroid. We demonstrate several methods for establishing spectral independence, including the correlation decay approach introduced by Weitz (2006) and the analytic approach introduced by Barvinok (2015) using zero-freeness of the associated complex-valued partition function. As a consequence of the techniques presented in this monograph we obtain fast mixing of the Gibbs sampler on general graphs in the so-called tree-uniqueness region, polynomial-time mixing on general graphs at the critical point for the uniqueness threshold, and polynomial-time mixing on random regular graphs beyond the uniqueness threshold. We also utilize the local-to-global framework of spectral independence and the Trickle-Down theorem of Oppenheim (2018) to prove fast mixing of the bases-exchange walk for generating a random basis of an arbitrary matroid, which was established by Anari, Liu, Oveis Gharan and Vinzant (2019).
Our focus in this monograph is on the analysis of the spectral gap of the associated Markov chains from a functional analysis perspective; we present proofs of the associated local-to-global theorems and the Trickle-Down Theorem from this same Markov chain perspective. The monograph is self-contained and aims to present the proofs in a unified fashion.
@@ 1 Introduction
This monograph aims to present the new technique of spectral independence together with its implications on the mixing time of the single-site update Markov chain known as the Glauber dynamics. Spectral independence is a simple, yet seemingly quite powerful technique to obtain optimal upper bounds on the convergence rate of Markov chains to its stationary distribution. We will introduce the spectral independence technique, explain the basic properties of the associated influence matrix, and discuss connections to simplicial complexes. We then prove that spectral independence implies fast mixing of the Glauber dynamics.
We also present the prominent result of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19] which shows fast mixing of a Markov chain, known as the bases-exchange walk, for randomly sampling bases of a matroid. This utilizes the local-to-global theorems presented for spectral independence and the Trickle-Down Theorem of Oppenheim [Opp18] to bound the spectral gap of the associated local walks.
Our aim in this work is to present the analyses from a Markov chain perspective, and to be as self-contained as possible.
@@ 1.1 Setting
Our setting is high-dimensional probability distributions defined by a graphical representation; this setting corresponds to the equilibrium distribution for spin systems in statistical physics and undirected graphical models in machine learning. Let $G=(V,E)$ be a graph on $n=|V|$ vertices. Consider a distribution $\mu$ on a subset of $\{0,1\}^{V}$. Let $\Omega\subseteq\{0,1\}^{V}$ be the support of $\mu$, that is,
$\Omega=\left\{\sigma\in\{0,1\}^{V}:\mu(\sigma)>0\right\}.$
We refer to the above setting as a binary system since each vertex has an assignment in $\{0,1\}$. Many of the results in these notes generalize to multispin systems which are distributions defined on $\{1,\dots,q\}^{V}$ for an integer $q\geq 2$; we will comment on the generalization to $q\geq 3$ spins at the end of this monograph in Section 12.2, and outside that section of the text we only consider the binary $q=2$ case.
An example to keep in mind is the so-called hard-core model that we now define.
@@ 1.2 Running example: Hard-core Model
The hard-core model is defined by a graph $G=(V,E)$ and an activity $\lambda>0$. The motivation of the hard-core model in statistical physics is as an idealized model of a gas where $\lambda$ corresponds to the fugacity of the gas (the fugactiy is related to the density). Configurations of the model are defined on independent sets of $G$: an independent set is a subset $I\subset V$ of vertices which does not contain an edge, which means that for all $\{u,v\}\in E$ either $u\notin I$ and/or $v\notin I$. Let $\Omega=\Omega_{G}$ denote the collection of independent sets of $G$ (regardless of their sizes).
For an independent set $\sigma\in\Omega$, we can view $\sigma$ as an $n$-dimensional indicator vector in $\{0,1\}^{n}$ where the $i$-th coordinate is assigned 1 if the $i$-th vertex is in $\sigma$ and is assigned 0 otherwise. Thus,
$\Omega=\Omega_{G}=\{\sigma\in\{0,1\}^{V}:\forall\{v,w\}\in E,\sigma(v)+\sigma(w)\leq 1\}.$
For an independent set $\sigma$, when a vertex $v\in\sigma$ (corresponding to $\sigma(v)=1$) then we often refer to $v$ as occupied, whereas for $w\notin\sigma$ (corresponding to $\sigma(w)=0$) we refer to $w$ as unoccupied. Each independent set $\sigma\in\Omega$ is assigned the following weight:
@@
where $|\sigma|$ is the number of vertices in the independent set $\sigma$, in other words, $|\sigma|=|\{v\in V:\sigma(v)=1\}|$. The Gibbs distribution $\mu=\mu_{G,\lambda}$ is the probability distribution on independent sets that is proportional to the weights. Specifically, for $\sigma\in\Omega$,
@@
where the normalizing factor $Z=Z_{G,\lambda}=\sum_{\eta\in\Omega}\lambda^{|\eta|}$ is known as the partition function.
1.3 Glauber dynamics/Gibbs sampler
The following Markov chain is a simple single-site update Markov chain designed to sample from the Gibbs distribution. The chain is known as the Glauber dynamics in the statistical physics community and as the Gibbs sampler in the machine learning community. We begin with the definition of the Glauber dynamics for the special case of the hard-core model (the reader new to the field can further consider the case $\lambda=1$ for simplicity).
@@ 1.3 Glauber dynamics/Gibbs sampler
    If $X^{\prime}\in\Omega$ (i.e., $X^{\prime}$ is a valid independent set) then set $X_{t+1}=X^{\prime}$, and otherwise set $X_{t+1}=X_{t}$.
Let $P$ denote the transition matrix of the Glauber dynamics. Note that $P$ is an $N\times N$ stochastic matrix where $N=|\Omega|$. The Glauber dynamics is irreducible and aperiodic and hence the Glauber dynamics is ergodic. To see that the dynamics is irreducible, let $\tau=\emptyset$ and notice that every state $\sigma\in\Omega$ can reach the empty set, i.e., $P^{t}(\sigma,\tau)>0$ for some $t\geq 0$, by successively removing vertices from $\sigma$, and similarly for every $\sigma^{\prime}\in\Omega$, $P^{t^{\prime}}(\tau,\sigma^{\prime})>0$ for some $t^{\prime}\geq 0$. Hence, $P^{s}(\sigma,\sigma^{\prime})>0$ for some $s\geq 0$ where this time $s=t+t^{\prime}$ depends on $\sigma,\sigma^{\prime}$ (though one can easily argue that there is a universal such time $s$). To see that the Glauber dynamics is aperiodic, notice that for every state $\sigma\in\Omega$ there is a so-called self-loop, namely $P(\sigma,\sigma)>0$, since the updated vertex $v$ can have $X^{\prime}(v)=X_{t}(v)$ with positive probability. Since the Glauber dynamics is ergodic, there is a unique stationary distribution $\mu$. Moreover, the Gibbs distribution $\mu$ is the unique stationary distribution, since it satisfies the detailed balance conditions, which state that for all pairs $\sigma,\sigma^{\prime}\in\Omega$, $\mu(\sigma)P(\sigma,\sigma^{\prime})=\mu(\sigma^{\prime})P(\sigma^{\prime},\sigma)$; hence, the Glauber dynamics is a reversible chain.
Before proceeding, we present the general definition of the Glauber dynamics for more general distributions. For a distribution $\mu$ on $\Omega\subset\{0,1\}^{V}$, the Glauber dynamics is defined as follows. From a state $X_{t}\in\Omega$,
@@
    Choose $X_{t+1}(v)$ from the conditional Gibbs distribution $\mu(\sigma(v)=\cdot\mid\sigma(w)=X_{t}(w)\mbox{ for all }w\neq v)$. In words, fix the spin/label at all vertices except $v$ and resample the spin/label at $v$ conditional on the fixed configuration on the rest of the vertices.
@@
The mixing time of the Glauber dynamics, denoted as $T_{\mathrm{mix}}(\varepsilon)$ for $\varepsilon>0$, is the minimum number of steps $t$ such that, for the worst initial state $X_{0}$, the distribution of $X_{t}$ is within total variation distance $\leq\varepsilon$ of the stationary distribution $\mu$. For a pair of distributions $\mu,\omega$ on $\Omega$, denote their total variation distance as follows:
$\|\mu-\omega\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\sigma\in\Omega}|\mu(\sigma)-\omega(\sigma)|=\max_{S\subset\Omega}\mu(S)-\omega(S).$
Then the mixing time is defined as:
$T_{\mathrm{mix}}(\varepsilon)=\min\{t:\forall X_{0}\in\Omega,\|\mu-P^{t}(X_{0},\cdot)\|_{\mathrm{TV}}\leq\varepsilon\}.$
Setting $\varepsilon=1/4$ we will refer to $T_{\mathrm{mix}}=T_{\mathrm{mix}}(1/4)$ as the mixing time. The choice of constant $1/4$ is somewhat arbitrary as one can “boost” the mixing time to obtain, for any $\varepsilon>0$, $T_{\mathrm{mix}}(\varepsilon)\leq T_{\mathrm{mix}}(1/4)\times\lceil\log_{2}(1/\varepsilon)\rceil$. Oftentimes, the constant $1/(2e)$ is used in place of $1/4$ in the definition of the mixing time and then the corresponding boosting has $\ln(1/\varepsilon)$ in place of $\log_{2}(1/\varepsilon)$, see [LP17, Section 4.5].
@@
We say the mixing time is optimal if $T_{\mathrm{mix}}=O(n\log{n})$ as this matches the lower bound established by Hayes and Sinclair [HS07] for any graph of constant maximum degree $\Delta$. Note an upper bound of $T_{\mathrm{mix}}=O(n\log{n})$ yields a more general bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log{n}\log(1/\varepsilon))$ via the above boosting result. However, the proofs of optimal mixing in these notes will yield a stronger upper bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$, and hence we will state this improved bound when relevant.
@@
In contrast to the mixing time, a weaker notion of convergence is the relaxation time. For $N=|\Omega|$, let $\lambda_{1}=1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-1$ denote the eigenvalues of the transition matrix $P$ of the Glauber dynamics (recall, the Glauber dynamics is reversible, i.e., satisfies the detailed balance conditions). Note, $\lambda_{1}$ corresponds to the principal eigenvector which is the stationary distribution $\mu$. Moreover, $\lambda_{2}<1$ since the Glauber dynamics is an ergodic Markov chain and hence the stationary distribution is unique. In addition, $\lambda_{N}>-1$ since the Glauber dynamics is reversible and aperiodic.
The relaxation time is defined as $T_{\mathrm{relax}}:=(1-\lambda_{*})^{-1}$ where $\lambda_{*}=\max\{\lambda_{2},|\lambda_{N}|\}$. The quantity $1-\lambda_{*}$ is known as the absolute spectral gap, see Section 3.4 for further discussion. The relaxation time measures the decay rate of the variance of the chain with respect to the stationary distribution; this will be formally presented in Section 3. The optimal relaxation time for the Glauber dynamics is $O(n)$, whereas the optimal mixing time is $O(n\log{n})$. An optimal mixing time of the form $O(n\log(n/\varepsilon))$ implies an optimal relaxation time (see Remark 9.30 in Section 9.5), but the reverse implication does not necessarily hold (see Section 3.7 for related results).
One of the primary applications of an (approximate) sampling algorithm is for the following problem of estimating the partition function. Consider the algorithmic problem of obtaining a Fully Polynomial-time Randomized Approximation Scheme ($\mathsf{FPRAS}$) for estimating the partition function $Z_{G,\lambda}$ of the hard-core model for a given $G=(V,E)$ and $\lambda\geq 0$; more precisely, our goal is to construct an algorithm which, given as input a graph $G=(V,E)$ and parameters $\lambda^{*}\geq 0$ and $\delta>0$, produces an estimate $\mathsf{EST}$ where:
$\mathrm{Pr}\left((1-\delta)Z_{G,\lambda}\leq\mathsf{EST}\leq(1+\delta)Z_{G,\lambda}\right)\geq 3/4,$
and runs in time $\mathrm{poly}(n,1/\delta)$ where $n=|V|$. (The success probability of $3/4$ can be boosted to $\geq 1-\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0$ with an additional $\log(1/\varepsilon^{\prime})$ factor in the running time.) Given a parameter $\lambda^{*}$, if for all $\lambda\leq\lambda^{*}$ we have an optimal mixing time of $O(n\log(n/\varepsilon))$ for the Glauber dynamics, then there is an $\mathsf{FPRAS}$ with running time $O((n/\delta)^{2}\log(n)\log(n/\delta))$ for estimating the partition function $Z_{G,\lambda^{*}}$ at $\lambda^{*}$, see [ŠVV09, Hub15, Kol18]. In fact, for the purposes of an approximate counting algorithm, it suffices to have a bound on the relaxation time (instead of the mixing time) of an appropriate sampling algorithm. In particular, if we have a relaxation time of $O(n)$ for all $\lambda\leq\lambda^{*}$ then we obtain an $\mathsf{FPRAS}$ with running time of $O^{*}((n/\delta)^{2})$ where the $O^{*}()$ notation hides logarithmic factors, see [ŠVV09, Section 7]. Moreover, we can estimate other quantities of interest, such as the number of independent sets of a particular size, see [HK25, DP23].
Further properties and fundamental definitions regarding Markov chains are presented in Section 3. We lower bound the spectral gap of the Glauber dynamics (and hence upper bound the relaxation time) using a functional analytic approach to analyze the decay rate of the variance of the Markov chain with respect to the stationary distribution; consequently, in Section 3.3 we introduce the variance functional and the Dirichlet form which is a measure of the local variation of a Markov chain. In Section 3.5 we present the notion of approximate tensorization of variance, which is a useful technique for lower bounding the spectral gap of the Glauber dynamics. The relations between the relaxation time and the mixing time of a Markov chain are established in Section 3.7.
1.4 Key Definitions: Spectral Independence and the Influence Matrix
@@
Spectral independence was introduced by Anari, Liu, and Oveis Gharan [ALO20]. It is defined by an $n\times n$ influence matrix which captures the pairwise influence or correlations between pairs of vertices.
@@
Let $G=(V,E)$ be a graph where $V=\{1,\dots,n\}$, and $\mu$ be a distribution on $\Omega$ where $\Omega\subseteq\{0,1\}^{V}$. Let $\Psi$ be the following real-valued $n\times n$ matrix; we refer to $\Psi$ as the influence matrix. For $1\leq i,j\leq n$, let
$\Psi_{\mu}(i,j)=\Psi(i\rightarrow j)$ $:=\mu\left(\sigma(j)=1\mid\sigma(i)=1\right)-\mu\left(\sigma(j)=1\mid\sigma(i)=0\right)$ (1)
$=\mathrm{Pr}_{\sigma\sim\mu}\left[\sigma(j)=1\mid\sigma(i)=1\right]-\mathrm{Pr}_{\sigma\sim\mu}\left[\sigma(j)=1\mid\sigma(i)=0\right].$
Note, when $i=j$ then $\Psi(i,i)=1$ and hence the diagonal entries of $\Psi$ are defined to be $1$ (see Remark 1.3 below for further discussion regarding an alternative definition of the diagonal entries).
A closely related notion is the modified influence matrix which we formally define in Section 2.5. The modified influence matrix differs from Eq. 1 in that the last term changes from $\mu\left(\sigma(j)=1\mid\sigma(i)=0\right)$ to $\mu\left(\sigma(j)=1\right)$. This leads to a weaker notion than spectral independence (we will later refer to the alternative notion as weak spectral independence); the connections between the influence matrix and modified influence matrix are discussed in further detail in Section 2.5.
We require the generalization of the above definitions to arbitrary “pinnings” which are a fixed assignment $\tau$ assigned to a subset $S$ of vertices. For a subset of vertices $S\subset[n]=\{1,\dots,n\}$, a pinning on $S$ is an assignment $\tau:S\rightarrow\{0,1\}$. Let $\Omega_{\tau}$ denote the assignments on all of $V$ which are consistent with $\tau$; more formally, for $S\subset V$ and a pinning $\tau:S\rightarrow\{0,1\}$, let
$\Omega_{\tau}:=\{\sigma\in\Omega:\sigma(S)=\tau(S)\},$
where $\sigma(S)$ is the restriction of $\sigma$ to $S$, and hence $\sigma(S)=\tau(S)$ iff $\sigma(v)=\tau(v)$ for all $v\in S$.
We only consider pinnings $\tau$ that can be extended to a valid configuration $\sigma$ on the entire vertex set $V$. Hence, we say a pinning $\tau$ is valid if $\Omega_{\tau}\neq\emptyset$. Let $\mathcal{P}$ denote the collection of all valid pinnings (see Section 4.1 for further introduction).
@@
For a pinning $\tau\in\mathcal{P}$, let $\mu_{\tau}$ denote the conditional Gibbs distribution, in other words, $\mu_{\tau}$ is the distribution $\mu$ conditional on the fixed assignment $\tau$ on $S$. For $\sigma\in\Omega_{\tau}$, we have
$\mu_{\tau}(\sigma)=\mu(\sigma\mid\sigma(S)=\tau)=\frac{\mu(\sigma(\overline{S})\cup\tau(S))}{\mu(\Omega_{\tau})},$
$\mbox{where}\ \ \ \mu(\Omega_{\tau})=\sum_{\eta\in\Omega_{\tau}}\mu(\eta)=\sum_{\begin{subarray}{c}\eta\in\Omega:\\ \eta(S)=\tau(S)\end{subarray}}\mu(\eta).$
For the case of the hard-core model on a graph $G=(V,E)$, pinnings correspond to induced subgraphs of $G$ in the following manner. In the hard-core model, for $S\subset V$, a valid pinning $\tau$ on $S$ is an independent set on $S$. Since $\tau$ is an independent set, an equivalent realization of the pinning $\tau$ is to consider the induced subgraph on $V^{\prime}\subset V$ where $V^{\prime}$ is formed by starting from $V$ and then proceeding as follows: for every vertex $v\in S$ where $\tau(v)=0$ then remove $v$ (this is because fixing $\tau(v)=0$ means $v$ is fixed to be outside of the independent set), and for every vertex $v\in S$ where $\tau(v)=1$ then remove $v$ and all $w\in N(v)$ (since $\tau(v)=1$ then $v$ is fixed to be in the independent set and hence all neighbors of $v$ are not in the independent set); let $V^{\prime}$ be the remaining vertices. Let $H$ be the induced subgraph on $V^{\prime}$. The distribution $\mu_{H}$ (which is the Gibbs distribution on $H$ without any pinning) is identical to the distribution $\mu_{\tau}$ (which is the conditional Gibbs distribution on $G$ with pinning $\tau$). Hence, for the case of the hard-core model, the collection of valid pinnings for $G$ can be realized by considering all induced subgraphs of $G$ (without any pinnings).
@@
Returning to the general setting, for a pinning $\tau\in\mathcal{P}$ where $\tau:S\rightarrow\{0,1\}$ for $S\subset V$, let $T\subset V\setminus S$ denote the set of vertices which are “free” in the following sense:
[…]
@@
Note, a vertex $i\in V\setminus(S\cup T)$ is “frozen”, i.e., it can only attain one spin and hence we can fix the configuration on that vertex and remove the corresponding two rows/columns from the matrix $\Psi_{\tau}$; henceforth, we only define the influence matrix on the free vertices.
@@
$\Psi_{\mu_{\tau}}(i\rightarrow j)$ $:=\mu_{\tau}\left(\sigma(j)=1\mid\sigma(i)=1\right)-\mu_{\tau}\left(\sigma(j)=1\mid\sigma(i)=0\right)$ (2)
$=\mu\left(\sigma(j)=1\mid\sigma(i)=1,\sigma(S)=\tau\right)-\mu\left(\sigma(j)=1\mid\sigma(i)=0,\sigma(S)=\tau\right).$
Since the empty pinning $\emptyset$ is always a valid pinning, thus $\Psi=\Psi_{\emptyset}$ is identical to the definition in Eq. 1. As in Eq. 1, observe that the diagonal entries of $\Psi_{\tau}$ are defined to be $1$, see Remark 1.3 for further discussion.
The matrix $\Psi$ can be asymmetric, and the entries of $\Psi$ can be positive or negative. Nevertheless, all of the eigenvalues of $\Psi$ are real and non-negative, this is established in Lemma 2.2 in Section 2.1. Intuitively, the fact that $\Psi$ has real non-negative eigenvalues is because it is closely related to the covariance matrix $\mathrm{Cov}$. In particular, the influence matrix $\Psi$ is equal to a positive diagonal matrix times the covariance matrix $\mathrm{Cov}$, and then the real non-negative eigenvalues of $\Psi$ follows from the positive semidefiniteness of the covariance matrix.
Since all eigenvalues of $\Psi$ are real valued we can denote the maximum eigenvalue by $\lambda_{\max}(\Psi)$. Now we can define the spectral independence condition as follows.
@@
For $\eta\geq 0$, we say that $\mu$ is $\eta$-spectrally independent if for all pinnings $\tau\in\mathcal{P}$,
$\lambda_{\max}(\Psi_{\mu_{\tau}})\leq 1+\eta.$ (3)
Note that the spectral independence condition only depends on the distribution $\mu$, and it has no explicit dependence on the Glauber dynamics. The definition was extended to non-binary spin spaces, such as the Potts model and colorings, in [FGYZ21, CGŠV21] (see also [CLV21a, BCC+22] for a general formulation); we present these generalizations and discuss the related results in Section 12.2.
When $\mu$ is a product distribution then $\eta=0$. Our goal is to show that $\eta$ is constant, and hence in some sense it is not far from a product distribution.